Two angles are complementary. The sum of the measure of the first angle and half the second angle is 70 degrees. Find the measures of the angle.
Answer by checkley77(12844)

x+y=90
x+y/2=70 now subtract
---------------
y-y/2=20
y/2=20
y=2*20
y=40 ans.
x+40=90
x=90-40
x=50 ans.
Proof:
50+40/2=70
50+20=70